It says to write the ratio using fractional notation. Do not simplify.
3 to 5 1/7. I am not sure how to write mixed fractions, but the 5 and 1/7 and a mixed number.
Thanks

Answer by checkley79(3341)

3/(5 1/7)=3/[(5*7+1)/7]=3/[(35+1)/7]=3/(36/7)=3*7/36=21/36=7/12 ans.